# Enhancing Speech Large Language Models with Prompt-Aware Mixture of Audio Encoders

### v1

@@ A.1 Details of Models and Datasets
Best-1 4.30 9.46 16.27 41.4% 32.96 15.04 13.13 24.04
2 $\surd$ $\surd$ - - 5.07 09.50 13.59 49.5% 35.47 16.16 9.38 33.71 06.00
2 $\surd$ - $\surd$ - 3.82 07.55 13.51 38.4% 35.43 15.55 8.29 29.79 05.83
2 $\surd$ - - $\surd$ 4.37 09.70 14.79 43.2% 35.33 16.62 9.62 31.72 06.50
2 - $\surd$ $\surd$ - 3.17 06.76 19.60 61.2% 31.70 14.89 9.84 35.93 05.83
@@
Best-2 3.17 06.76 13.51 61.2% 35.47 16.62 9.85 36.65
3 $\surd$ $\surd$ $\surd$ - 3.65 07.07 12.79 42.8% 35.47 15.70 7.83 31.32 04.00
3 $\surd$ $\surd$ - $\surd$ 4.42 10.26 13.47 47.4% 35.32 16.62 9.38 33.11 06.50
3 $\surd$ - $\surd$ $\surd$ 4.58 06.99 15.11 59.1% 35.33 16.62 8.89 37.02 05.50
@@
Best-3 3.09 06.64 12.79 59.1% 35.47 16.62 9.24 31.45
4 $\surd$ $\surd$ $\surd$ $\surd$ 3.94 07.28 14.06 57.3% 35.37 16.79 8.43 36.49 04.00
@@
A.2 Details of Baseline Implement
In our work, we compare our method against two types of baselines. The first baseline consists of models using a single encoder, while the second baseline involves fusing multiple audio encoders either based on previous work Hu et al. (2024); Tang et al. (2024) or through an averaging operation.
For the single encoder baseline, we train the model using the same settings as in our method. For the second baseline, we train the model using the open-source codebases from WavLLM and SALMONN, we integrate the adapter components from these repositories into our code and train the baseline model using our training data, employing the same pre-trained audio encoders and LLMs as in our method. During training, we applied the same hyperparameters as our method. Since we use different encoders and LLMs compared to the baselines, we adjust the dimensions of the adapter to match the specific audio encoder and LLM we used while maintaining other dimensions independent of the audio encoders and LLM unchanged. Notably, we trained the SALMONN with query length=32 (as training with the original setting query length=1 failed) to ensure comparable performance with the other baseline methods.
A.3 All result

### v2

@@ A.10 All Detailed Results
Best-1 4.30 09.46 16.27 41.4% 32.96 15.04 13.13 24.04
2 $\surd$ $\surd$ - - 5.07 09.50 13.59 49.5% 35.47 16.16 09.38 33.71 06.00
2 $\surd$ - $\surd$ - 3.82 07.55 13.51 38.4% 35.43 15.55 08.29 29.79 05.83
2 $\surd$ - - $\surd$ 4.37 09.70 14.79 43.2% 35.33 16.62 09.62 31.72 06.50
2 - $\surd$ $\surd$ - 3.17 06.76 19.60 61.2% 31.70 14.89 09.84 35.93 05.83
@@
Best-2 3.17 06.76 13.51 61.2% 35.47 16.62 09.85 36.65
3 $\surd$ $\surd$ $\surd$ - 3.65 07.07 12.79 42.8% 35.47 15.70 07.83 31.32 04.00
3 $\surd$ $\surd$ - $\surd$ 4.42 10.26 13.47 47.4% 35.32 16.62 09.38 33.11 06.50
3 $\surd$ - $\surd$ $\surd$ 4.58 06.99 15.11 59.1% 35.33 16.62 08.89 37.02 05.50
@@
Best-3 3.09 06.64 12.79 59.1% 35.47 16.62 09.24 31.45
4 $\surd$ $\surd$ $\surd$ $\surd$ 3.94 07.28 14.06 57.3% 35.37 16.79 08.43 36.49 04.00
@@
A.11 Prompts in Training and Inference Stage
In practice, large speech-language models typically address downstream tasks using prompts that are semantically explicit but textually diverse, as illustrated in Table 14. Unlike classical MoE routing methods that rely on hidden states, or other approaches such as predefined task labels (e.g., “Task ID = 3”) Kudugunta et al. (2021) or random routing Zuo et al. (2021), our model leverages natural language prompts to route inputs to the appropriate expert.
These prompts vary in phrasing but convey the same task intent. We utilize the semantic understanding capabilities of a pretrained LLM to extract task information from these prompts, rather than depending on fixed task labels. As an example, in our ASR task, we trained the routing module using 50 diverse prompts (Table 15). During evaluation, the model was tested on 200 prompts, 150 of which were unseen during training (Table 16). Our prompt-based router, powered by the LLM, achieved 100% expert selection accuracy, demonstrating strong generalization to previously unseen prompts.
Convert the audio speech into a text transcript.
Write an accurate version of the audio content in text.
Capture what is being said in this audio as text.
Table 14: Example of common prompts used in ASR tasks.
Convert the audio speech into a text transcript.
Listen to the following audio and create a corresponding text transcript.
Transform the speech into a text document.
Listen to the audio and generate a text version of it.
Create a text version based on the audio speech provided.
Table 15: Example of prompts used during training stage.
Transcribe what is said in the audio into written text. (unseen)
Extract the words from the audio and write them down. (unseen)
Produce a typed version of the audio’s spoken content. (unseen)
Capture what is being said in this audio as text. (unseen)
Hear the provided audio and provide a written text version.
Table 16: Example of prompts used during inference stage.
A.12 Efficient
The low decoding efficiency of LLMs is due to repeated invocations of the whole decoder layer during autoregressive generation. Several methods have been proposed to address this issue, including speculative sampling Leviathan et al. (2023) and KV-cache compression Zhang et al. (2023). Recently, researchers have explored using LLMs as encoders Luo et al. (2025), thereby leveraging their knowledge while improving decoding efficiency. Additionally, enhancing the computational efficiency and performance of adapters in end-to-end speech models through optimized FFN dimension design Zheng et al. (2023) represents a viable solution.